Within each four hours worked, workers can take a total of 48 minutes in allowance (so allowance factor is based on workday). If the observed time is 6 minutes, and performance rating is 95%, what is the standard time (in minutes; give three significant digits after the decimal point)?
Answer #1

Standard time = Normal time * allowance factor

Where,

Normal time = observed time * performance rating

Observed time = 6 minutes and performance rating = 95%

Therefore,

Normal time = 6 minutes * 95% = 5.7 minutes

Within each four hours worked, workers can take a total of 48 minutes in allowance -

Allowance factor = 1+ (time taken as allowance/total time between allowance) = 1 + (48 / 4 *60)

= 1+ (48/ 240) = 1+ 0.2 or 120%

Therefore,

Standard time = 5.7 minutes *120% = 6.840 minutes